# Deeper Diffusion Models Amplify Bias

### v1

@@ Abstract
Despite the impressive performance of generative Diffusion Models (DMs), their internal working is still not well understood, which is potentially problematic. This paper focuses on exploring the important notion of bias-variance tradeoff in diffusion models. Providing a systematic foundation for this exploration, it establishes that at one extreme the diffusion models may amplify the inherent bias in the training data and, on the other, they may compromise the presumed privacy of the training samples. Our exploration aligns with the memorization-generalization understanding of the generative models, but it also expands further along this spectrum beyond “generalization”, revealing the risk of bias amplification in deeper models. Building on the insights, we also introduce a training-free method to improve output quality in text-to-image and image-to-image generation. By progressively encouraging temporary high variance in the generation process with partial bypassing of the mid-block’s contribution in the denoising process of DMs, our method consistently improves generative image quality with zero training cost. Our claims are validated both theoretically and empirically.
Figure 1: Deeper DMs tend to generate samples of majority classes more than shallow models. Unconditional DMs with different depths are trained on CelebA [20] that contains “blond male” group and “male” class as minorities (less than 1% and 41.68%, respectively). Starting from the same seeds on which shallow models generate samples of the minority, deeper models generate samples of the majority. (Left) Instead of “blond male”, deeper models generate “non-blond male” and “blond female”. (Right) Deeper models generate samples of the majority class “female”.
@@ 1 Introduction
Diffusion Models (DMs) show unprecedented performance in generating hyper-realistic and diverse synthetic images [5, 8, 22]. Although they have seen rapid empirical progress and widespread adoption in various applications [19, 12, 18, 30, 28], their underlying mechanisms and inference dynamics remain comparatively underexplored from a theoretical perspective. This growing gap between practical performance and foundational insight highlights the need for deeper analytical study which could accelerate the costly training process by diminishing the need for trial and error, and could possibly address bias and privacy concerns.
@@
In this work, we explore the traditional concept of bias-variance tradeoff [10] in the scope of DMs. Intuitively, our definition of this tradeoff in DMs is roughly analogous to the bias-variance tradeoff in k-Nearest Neighbors (k-NNs). We conceptualize the training of DMs as implicit memorization of the training data, as opposed to explicit storing of training samples in k-NNs. We show that higher levels of abstraction in the encoding process of DMs is equivalent to larger k values in k-NN. So, extremely low abstraction level leads to “high variance” and extensively high abstraction level results in “high bias”. We think of the denoising process of DMs as memory retrieval [11], which enables us to set a clear analogy between the way k-NNs and DMs treat the training data.
@@
Relevant to our exploration, the concepts of memorization and generalization are also previously studied in the context of DMs [32, 16, 17, 25]. Connecting our analysis to those studies, the established concept of “memorization” falls close to “high variance” in our work, while “generalization” falls somewhere between the two extremes of “high variance” and “high bias”. Consequently, our perspective enables more comprehensive insights, naturally revealing that “memorization” and “generalization” are not the ultimate extremes, issuing the caution about going far in the direction of “generalization” that can lead to “high bias” in the models.
Based on the memory-like interpretation of DMs and our bias-variance tradeoff explanation, we are also able to demonstrate a training-free method to enhance the visual quality of images generated by the DMs. The crux of our method lies in encouraging the exploitation of high-variance choices during inference by letting the generative process have more direct access to multiple local minima in the energy landscape. This can be achieved in a training-free manner by partially bypassing the mid-block in the denoiser. We demonstrate a consistent image quality improvement by our method.
Figure 2: Illustration of the key concepts. (A) Comparison of how Associative Memory (AM), k-NNs, and DMs store and attend to the training data. In AMs and k-NNs, data is stored explicitly, as opposed to implicit storing in DMs. Corrupted signals (noise) in DMs converge to attractors like AMs. However, in DMs there is no guarantee to converge to a memory, but the attractors are potentially superpositions of memories. Like k-NNs, DMs attend to a number of training samples during inference. (B) Top: For smoother energy landscapes, initial corrupted signals tend to attend to a larger number of memories (High Bias), which leads to amplification of the inherent dataset bias. We show actual bias amplification of absolute 12.82% in CelebA [20] model. This leads to more frequent generation of the majority class. E.g., “Non-Blond Male” and “Female” are more frequently generated for CelebA by deeper models, which naturally inherit higher bias as per our analysis. Bottom: For less smooth energy landscapes, initial corrupted signals are likely to attend to a smaller number of memories (High Variance). High variance restricts initial corrupted signals to converge to attractors corresponding to a small number of memories, leading to non-diverse generation, potentially revealing training data that is private. We show examples generated by the prompts “An actor" and “An actress". Starting from the same gaussian seed we gradually push SD V1.5 [23] to higher variance state during the inference. Both cases eventually converge to the same/similar person. (C) Our explanation of “bias-variance tradeoff” for DMs covers the contrasting concepts of “memorization” and “generalization”. However, it establishes a more complete perspective beyond “generalization”. It captures the risk of falling into “high bias” state by getting too far from “memorization”.
The key contributions of this paper are:
@@
  • We establish that the abstraction level of DM directly impacts bias-variance tradeoff, as extensively abstracted representation leads to exacerbation of bias, and conversely, low abstraction level can undermine the presumed privacy of DMs.
  • Leveraging insights from our analysis, we propose a training-free method to improve image generation quality for DMs. We extensively evaluate our findings and proposed approach.
@@ 2 Diffusion Models as Implicit Memory Modules
To facilitate our “bias-variance tradeoff” explanation, first we need to conceptualize the generative process of DMs as a “memory recall” task. As confirmed in [2], DMs can memorize specific training samples and output them during generation. They implicitly store data in latent representations, enabling lossy recall via iterative refinement. This perspective emphasizes the role of the model’s learned score function as implicitly encoding a memory retrieval mechanism.
To further explain the memory retrieval behavior, we employ the analogy between energy-based Associative Memories (AMs) [14] and DMs, inspired by [11]. Under this perspective, we view the latent space of DMs as an energy landscape, in which high energy corresponds to noisy signals and low energy represents images resembling training data distribution. The denoising process starts with a high energy corrupted signal, and by following the direction of negative energy gradients, it eventually ends up in an attractor, i.e., a local minimum. In DMs, an attractor can be a “memory” or a “superposition of memories” (see Figure 2 (A)). The learning phase of DMs can be viewed analogous to traditional memory-based algorithms, such as k-NN, in the sense that they both store training data representations. However, it differs to k-NN in explicit-implicit dichotomy. In our interpretation of the training process of DMs, training data is implicitly memorized by DMs, as opposed to the explicit memorization in k-NN. Hence, we treat DMs as “Implicit Memory Modules”.
@@ 3 Bias-Variance Tradeoff in Diffusion Models
Given the analogy between the implicit memorization in DMs and explicit memorization in k-NNs, we can gain an insight on bias-variance tradeoff in DMs. In k-NNs the value of the hyperparameter k determines this tradeoff. Higher value of k leads the query point to attend to a larger number of neighbors, which smooths the decision landscape and reduces sensitivity to individual data points. Similarly, in the context of DMs, an initial corrupted noise (comparable to a query in k-NN) attends to a number of implicit memories (comparable to explicit memories in k-NN) to generate a sample resembling the distribution of a class  [^1] We intentionally treat “concepts” as “classes” to relate our discussion to k-NNs. (comparable to categorization to a class in k-NN.). Analogous to the higher values of k in k-NNs, which lead to the smoothening of the decision landscape, energy landscape of DMs smoothens with higher levels of abstraction in the DM representations. Hence, we can expect bias-variance trade-off in DMs to be linked with the abstraction level of representations.
Let $D=\{x_{i}\}_{1}^{N}\subset\mathbb{R}^{d}$, where $D$ is training dataset and each $x_{i}$ is a training sample, and $q\in\mathbb{R}^{d}$ denote a query point (a corrupted signal). In k-NN, prediction is based on the average over the k closest training points to q, denoted as $\mathcal{N}_{k}(q)\subset D$. The k-NN prediction can be written as:
@@
where $\tau$ controls the rate of change, iteratively mapping $q$ to a nearby energy minimum. The smoothness of the energy landscape, defined by the number, sharpness, and separation of local minima, determines how many minima influence the trajectory of $q$. In a rugged landscape, the flow quickly collapses into a nearby basin, analogous to small-$k$ behavior in k-NN, where only the most immediate neighbors contribute. In contrast, a smoothed energy landscape causes the flow to be influenced by multiple nearby minima, resembling a k-NN with large k aggregating information from many neighbors. Thus, $\tau$ has the same role in DMs as $k$ in k-NNs. A comprehensive set of analogies between DMs and k-NNs is given in Table 1.
Table 1: Analogies between k-NNs and DMs.
@@
Inference process attends to nearest neighbors converges to a minimum corresponding to nearby memories
Smoothing factor k directly defined as a hyperparameter $\tau$ defined by abstraction level of representation
Larger smoothing factor smoothens the decision boundary smoothens the energy landscape
Inference goal classify the query refine the corrupted signal to an image resembling a class in training data
@@ 3.1 Implications of Latent Space Abstraction Level
Here we examine how the energy landscape corresponding to DMs’ learned memories get affected by the abstraction level of their latent space.
As formally depicted in Lemma 1, increasing the level of abstraction in latent representations induces smoother energy landscapes. To support this claim, we analyze the local and global smoothness of the energy function by examining two key quantities: the Hessian norm, which captures local curvature, and the Lipschitz constant of the gradient field, which reflects how rapidly gradients can vary across the space. We show that both the Hessian norm and the Lipschitz constant decrease with higher levels of abstraction.
Lemma 1 (Higher abstraction in representations induces smoother energy landscapes).
@@
The complete proof of Lemma 1 is given in the Appendix. As a result of smoothing energy landscapes, some of the nearby minima will be merged in to a wider yet shallower minimum basin, which is defined in Definition 1.
@@
Let $E:\mathbb{R}^{d}\to\mathbb{R}$ be an energy function corresponding to a latent space. If the energy landscape becomes smoother, then a set of nearby local minima $\{x_{i}^{*}\}_{i=1}^{n}$ of $E$ may merge into a single, wider and shallower local minimum $\tilde{x}^{*}$ of the smoothed energy function $\tilde{E}$. We refer to such a minimum $\tilde{x}^{*}$ as a Merged Local Minimum.
The local minima that form a merged local minimum can be potentially from different classes. The question that arises here is what happens if the high-energy corrupted signal ends up in such a new merged minimum. In other words, what is the probability of generating each of the classes residing in the merged minimum? Answering this, we show in Lemma 2 that the generation gets heavily biased in favor of the majority class.
Lemma 2 (Merged local minima are biased representatives).
@@
$H_{C}\{x_{i}^{*}\}_{i=1}^{n}):=\left\{\sum_{i=1}^{n}\alpha_{i}x_{i}^{*}\mid\alpha_{i}\geq 0,\sum_{i=1}^{n}\alpha_{i}=1\right\}.$ (6)
@@
The complete proof of Lemma 2 is presented in the Appendix, and an intuitive illustration of the consequences of biased representatives is presented in Figure 3, which shows five energy landscapes with different proportion of classes. The stronger the initial bias, the greater the bias amplification induced by representation abstraction.
Figure 3: Five different energy landscapes of size 32x64 containing minima for two classes: “blue” and “yellow” with corresponding proportions of 50%-50%, 60%-40%, 70%-30%, 80%-20%, and 90%-10% are shown (left column). Each of the landscapes gets gradually more abstract by getting local average over classes using a 2x2 window to produce 16x32, 8x16, 4x8, and 2x4 representations. The proportion of minina for the 50%-50% case remains almost the same for every level of abstraction (see first row). However, for biased cases the bias gets amplified, and the amplification rate gets exacerbated for higher levels of initial bias (see rows 2 to 5). The percentage of the “blue” class is given above each landscape.

### v2

@@ Abstract
Despite the remarkable performance of generative Diffusion Models (DMs), their internal working is still not well understood, which is potentially problematic. This paper focuses on exploring the important notion of bias-variance tradeoff in diffusion models. Providing a systematic foundation for this exploration, it establishes that at one extreme, the diffusion models may amplify the inherent bias in the training data, and on the other, they may compromise the presumed privacy of the training samples. Our exploration aligns with the memorization-generalization understanding of the generative models, but it also expands further along this spectrum beyond “generalization”, revealing the risk of bias amplification in deeper models. Our claims are validated both theoretically and empirically.
@@ 1 Introduction
Diffusion Models (DMs) show unprecedented capabilities in generating hyper-realistic and diverse synthetic images (Dhariwal & Nichol, 2021; Esser et al., 2021; Podell et al., 2023). Although they have seen rapid empirical progress and widespread adoption in various applications (Liu et al., 2023; Huang et al., 2023; Li et al., 2022; Wu et al., 2023; Watson et al., 2023), their underlying mechanisms and inference dynamics remain comparatively underexplored from a theoretical perspective. This growing gap between practical performance and foundational insight highlights the need for deeper analytical study, which could accelerate the costly training process by diminishing the need for trial and error, and possibly address bias and privacy concerns.
Figure 1: Deeper Diffusion Models (DMs) tend to generate samples of majority classes more compared to shallower models, while shallower models tend to give away information about their training data. Here, unconditional DMs with different depths are trained on CelebA (Liu et al., 2015) that contains “blond male” group and “male” class as minorities (less than 1% and 41.68%, respectively). High Bias: Starting from the same seeds on which shallow models generate samples of the minority, deeper models generate samples of the majority. Instead of “blond male”, deeper models generate “non-blond male” and “blond female”. Also, deeper models generate samples of the majority class “female” instead of “male” . High Variance: Shallower models are prone to generate images similar to the training dataset samples. The generated image of the shallow model resembles its nearest neighbor much more compared to the image generated by the deep model.
@@
In this work, we explore the traditional concept of bias-variance tradeoff (Geman et al., 1992) in the scope of DMs. Intuitively, our definition of this tradeoff in DMs is roughly analogous to the bias-variance tradeoff in k-Nearest Neighbors (k-NNs). We conceptualize the training of DMs as implicit memorization of the training data, as opposed to explicit storing of training samples in k-NNs. We show that higher levels of abstraction in the encoding process of DMs are equivalent to larger k values in k-NN. So, extremely low abstraction levels lead to “high variance”, while excessively high abstraction levels result in “high bias” (see Figure˜1). We conceptualize the denoising process of DMs as memory retrieval (Hoover et al., 2023), which enables us to set a clear analogy between the way k-NNs and DMs treat training data.
@@
Relevant to our exploration, the concepts of memorization and generalization are also previously studied in the context of DMs (Yoon et al., 2023; Li et al., 2023; 2024; Vastola, 2025). Connecting our analysis to those studies, the established concept of “memorization” falls close to “high variance” as per our analysis, while “generalization” falls somewhere between the two extremes of “high variance” and “high bias”. Consequently, our perspective enables more comprehensive insights, naturally revealing that “memorization” and “generalization” are not the ultimate extremes, issuing the caution about going far in the direction of “generalization” that can lead to “high bias” in the models.
Our key contributions are summarized below:
@@
  • We establish that the abstraction level of DM directly impacts bias-variance tradeoff, as extensively abstracted representations lead to exacerbation of bias, while conversely, low abstraction levels can undermine the presumed privacy of DMs.
Figure 2: Illustration of the key concepts. (A) Comparison of how Associative Memory (AM), k-NNs, and DMs store and attend to the training data. In AMs and k-NNs, data is stored explicitly, as opposed to implicit storing in DMs. Corrupted signals (noise) in DMs converge to attractors like AMs. However, in DMs, there is no guarantee to converge to a memory, but the attractors are potentially superpositions of memories. Like k-NNs, DMs attend to a number of training samples during inference. (B) Top: For smoother energy landscapes, initial corrupted signals tend to attend to a larger number of memories (High Bias), which leads to amplification of the inherent dataset bias. We show actual bias amplification of absolute 12.82% in CelebA (Liu et al., 2015) model. This leads to more frequent generation of the majority class. E.g., “Non-Blond Male” and “Female” are more frequently generated for CelebA by deeper models, which naturally inherit higher bias as per our analysis. Bottom: For less smooth energy landscapes, initial corrupted signals are likely to attend to a smaller number of memories (High Variance). High variance restricts initial corrupted signals to converge to attractors corresponding to a small number of memories, leading to non-diverse generation, potentially revealing training data that is private. We show examples generated by the prompts “An actor" and “An actress". Starting from the same gaussian seed we gradually push SD V1.5 (Rombach et al., 2022) to a higher variance state during the inference. Both cases eventually converge to the same/similar person. (C) Our explanation of “bias-variance tradeoff” for DMs covers the contrasting concepts of “memorization” and “generalization”. However, it establishes a more complete perspective beyond “generalization”. It captures the risk of falling into “high bias” state by getting too far from “memorization”.
@@ 2 Diffusion Models as Implicit Memory Modules
To facilitate our “bias-variance tradeoff” explanation, first we need to conceptualize the generative process of DMs as a “memory recall” task. As confirmed in (Carlini et al., 2023), DMs can memorize specific training samples and output them during generation. They implicitly store data in latent representations, enabling lossy recall via iterative refinement. This perspective emphasizes the role of the model’s learned score function as implicitly encoding a memory retrieval mechanism.
To further explain the memory retrieval behavior, we employ the analogy between energy-based Associative Memories (AMs) (Krotov & Hopfield, 2021) and DMs, inspired by (Hoover et al., 2023). Under this perspective, we view the latent space of DMs as an energy landscape, in which high energy corresponds to noisy signals and low energy represents images resembling the training data distribution. The denoising process starts with a high energy corrupted signal, and by following the direction of negative energy gradients, it eventually ends up in an attractor, i.e., a local minimum. In DMs, an attractor can be a “memory” or a “superposition of memories” (see Figure˜2 (A)). The learning phase of DMs can be viewed analogously to traditional memory-based algorithms, such as k-NN, in the sense that they both store training data representations. However, it differs to k-NN in explicit-implicit dichotomy. In our interpretation of the training process of DMs, training data is implicitly memorized by DMs, as opposed to the explicit memorization in k-NN. Hence, we treat DMs as “Implicit Memory Modules”.
@@ 3 Bias-Variance Tradeoff in Diffusion Models
Given the analogy between the implicit memorization in DMs and explicit memorization in k-NNs, we can gain an insight on the bias-variance tradeoff in DMs. In k-NNs, the value of the hyperparameter k determines this tradeoff. Higher value of k leads the query point to attend to a larger number of neighbors, which smooths the decision landscape and reduces sensitivity to individual data points. Similarly, in the context of DMs, an initial corrupted noise (comparable to a query in k-NN) attends to a number of implicit memories (comparable to explicit memories in k-NN) to generate a sample resembling the distribution of a class  [^1] We intentionally treat “concepts” as “classes” to relate our discussion to k-NNs. (comparable to categorization to a class in k-NN.). Analogous to the higher values of k in k-NNs, which lead to the smoothening of the decision landscape, the energy landscape of DMs smoothens with higher levels of abstraction in the DM representations (see Figure˜2). Hence, we can expect the bias-variance trade-off in DMs to be linked with the abstraction level of representations.
Consider a dataset $D=\{x_{i},y_{i}\}_{i=1}^{N}\subset\mathbb{R}^{d}$, where each $x_{i}$ is a training sample and $y_{i}$ is its corresponding label, and $Q\in\mathbb{R}^{d}$ denote a query point (a corrupted signal). In k-NN, prediction is based on the average over the k closest training points to $Q$, denoted as $\mathcal{N}_{k}(Q)\subset D$. The k-NN prediction can be written as:
@@
where $\tau$ controls the rate of change, iteratively mapping $Q$ to a nearby energy minimum. The smoothness of the energy landscape, defined by the number, sharpness, and separation of local minima, determines how many minima influence the trajectory of $Q$. In a rugged landscape, the flow quickly collapses into a nearby basin, analogous to small-$k$ behavior in k-NN, where only the most immediate neighbors contribute. In contrast, a smoothed energy landscape causes the flow to be influenced by multiple nearby minima, resembling a k-NN with large k aggregating information from many neighbors. Thus, $\tau$ has the same role in DMs as $k$ in k-NNs. A comprehensive set of analogies between DMs and k-NNs is given in Table˜1.
Table 1: Analogies between k-NNs and Diffusion Models (DMs).
@@
Inference process attends to nearest neighbors converges to a minimum in nearby memories
Smoothing factor k defined as a hyperparameter $\tau$ defined by abstraction level of representation
Larger smoothing factor smoothens decision boundaries smoothens the energy landscape
Inference goal classify the query refine corrupted signal to resemble a class
Our definition of “bias-variance tradeoff” aligns with its traditional counterpart, specifically with k-NN classification. The bias and the variance of the model $\hat{f}(x)$ is defined by
$bias(\hat{f}(x_{0}))=\mathbb{E}[\hat{f}(x_{0})]-f(x_{0})],$ (4)
and
$variance(\hat{f}(x_{0}))=\mathbb{E}\left[(\hat{f}(x_{0})-\mathbb{E}[\hat{f}(x_{0})])^{2}\right],$ (5)
where specifically in our definition of bias-variance in diffusion models, $f(x_{0})$ is the class of the training image that is encoded to $x_{0}$ and $\hat{f}(x_{0})$ is the class of the generated image decoded from $x_{0}$.
@@ 3.1 Implications of Latent Space Abstraction Level
Here, we examine how the energy landscape corresponding to DMs’ learned memories gets affected by the abstraction level of their latent representation.
As formally depicted in Theorem˜1, increasing the level of abstraction in latent representations induces smoother energy landscapes. To support this claim, we analyze the local and global smoothness of the energy function by examining two key quantities: the Hessian norm, which captures local curvature, and the Lipschitz constant of the gradient field, which reflects how rapidly gradients can vary across the space. We show that both the Hessian norm and the Lipschitz constant decrease with higher levels of abstraction.
Theorem 1 (Higher abstraction in representations induces smoother energy landscapes).
@@
The complete proof of Theorem˜1 is given in the Appendix. As a result of smoothing energy landscapes, some of the nearby minima will be merged into a wider yet shallower minimum basin, which is defined in Definition˜1.
@@
Let $E:\mathbb{R}^{d}\to\mathbb{R}$ be an energy function corresponding to a latent representation. If the energy landscape becomes smoother, then a set of nearby local minima $\{x_{i}^{*}\}_{i=1}^{n}$ of $E$ may merge into a single, wider and shallower local minimum $\tilde{x}^{*}$ of the smoothed energy function $\tilde{E}$. Formally, for some small $\varepsilon>0$,
$\|\phi(x_{i}^{*})-\tilde{x}^{*}\|\leq\varepsilon,\quad\forall i=1,\dots,n,$
where $\phi$ is an abstraction (encoding) map. Hence, $\tilde{x}^{*}$ is a representative of the local minima in $E$ under the abstraction i.e. in $\tilde{E}$. We refer to such a minimum $\tilde{x}^{*}$ as a “Merged Local Minimum”.
The local minima that form a merged local minimum can potentially be from different classes. The question that arises here is what happens if the high-energy corrupted signal ends up in such a new merged minimum. In other words, what is the probability of generating each of the classes residing in the merged minimum? Answering this, we show in Theorem˜2 that the generation gets heavily biased in favor of the majority class.
Theorem 2 (Merged local minima are biased representatives).
@@
$H_{C}(\{x_{i}^{*}\}_{i=1}^{n}):=\left\{\sum_{i=1}^{n}\alpha_{i}x_{i}^{*}\mid\alpha_{i}\geq 0,\sum_{i=1}^{n}\alpha_{i}=1\right\}.$ (8)
@@
Proof.
Let $\tilde{x}^{*}$ be a merged local minimum basin formed by smoothing the energy landscape containing $n$ local minima, with the number of local minima corresponding to classes $A$ and $B$ being $p$ and $q$, respectively. The initial probability of convergence to classes $A$ and $B$ before smoothing is
$\mathbb{P}_{\text{init}}(A)=\frac{p}{n},\quad\mathbb{P}_{\text{init}}(B)=\frac{q}{n}.\vskip-1.42262pt$ (13)
The initial odds ratio of class $A$ to class $B$ is given by
$\lambda_{\text{init}}(A)=\frac{p}{q}.\vskip-5.69054pt$ (14)
After merging the minima, the generated image will belong to class $A$, class $B$, or a mixture of both. However, as long as classes $A$ and $B$ are distinguishable and the model is well-trained, the generation of a mixture of classes is not possible. Therefore, the generated image must either be from class $A$ or from class $B$.
To avoid generating a mixture, the model must select all exclusive semantic features from one class. Thus, to generate an image from class $A$, all exclusive features must be selected from class $A$, each with an odds ratio of $\frac{p}{q}$. Let $S$ be the number of exclusive semantic features. The odds ratio of generating an image from class $A$ to class $B$ after smoothing is
$\lambda_{\text{smooth}}(A)=\left(\frac{p}{q}\right)^{S}.\vskip-5.69054pt$ (15)
If $p>q$, for $S>1$ we have
$\lambda_{\text{smooth}}(A)=\left(\frac{p}{q}\right)^{S}>\frac{p}{q}=\lambda_{\text{init}}(A),\vskip-5.69054pt$ (16)
and for a sufficiently large $S$,
$\lambda_{\text{smooth}}(A)\gg\lambda_{\text{init}}(A).$ (17)
So,
$\mathbb{P}_{\text{smooth}}(A)\gg\mathbb{P}_{\text{smooth}}(B),$ (18)
where $\mathbb{P}_{\text{smooth}}(A)$ and $\mathbb{P}_{\text{smooth}}(B)$ are probabilities of convergence to classes A and B after smoothing, respectively.
Figure˜3 illustrates how biased representatives exacerbate bias. It shows five energy landscapes with different proportions of classes, getting gradually more abstract (left to right). Based on Theorem 2, we represent nearby classes as the majority class (random in case of a tie) in the more abstract representations. It can be seen that while class proportions remain almost identical in the balanced case, the majority class dominates with higher levels of abstraction in unbalanced cases. Also, the stronger the initial bias, the greater the bias amplification induced by representation abstraction.
Figure 3: (Illustration of bias amplification.) Five different energy landscapes containing minima for two classes: “red” and “blue” with corresponding proportions of 50%-50%, 60%-40%, 70%-30%, 80%-20%, and 90%-10% are shown (left column). Each of the landscapes gets gradually more abstract, and the class of the minima is decided based on Theorem 2 by getting a local average over classes using a 2x2 window (random in case of tie in the average). The proportion of minima for the 50%-50% case remains almost the same for every level of abstraction (see first row). However, for biased cases, the bias gets amplified, and the amplification rate gets exacerbated for higher levels of initial bias (see rows 2 to 5). The percentage of the “red” (majority) class is given above each landscape.